A solution is made by dissolving 75.0 grams glycerin C3H803 (92 g/mol), in 600.0 grams of water (18 g/mol) to make approximately 600 -mL solution. Calculate the following:


a. Vapor pressure of the solution at 25.0 Celsius (The vapor pressure of pure water is 23.8 torr at 25.0 Celsius).


b. Boiling point of the solution


c. Freezing point of the solution


d. Osmotic pressure at 25.0 Celsius


1
Expert's answer
2021-06-03T05:25:50-0400

a) psol = X solvent × Psolvent

75/92 = 0.82 moles glycerin

600/18 = 33.33 moles water

Total moles = 0.82 + 33.33 = 34.15

Mole fraction of water = 33.33/34.15

= 0.976

= 0.976 × 23.8 = 23.23 torr

b) 100 + 23.23 = 123.23°C

C) 0 -23.23 = -23.23°C